Assume you have a 6% 30-year mortgage for $100,000 with now 10 years to maturity (annual payments with exactly one year to the next payment). You are considering a refinance of the loan at 4% with a refinancing fee of $4,000.

What is the new payment, if you choose to refinance. Is refinancing with the new rate (4%) a good idea?
Answer #1

New payment = $7085.56

Original loan PV 100000
Rate Rate 6%
Term Nper 30
Number of payments done 20
Annual payment PMT $7,264.89
Principal paid off in 20 payments -46529.77
Balance on loan remaining 53470.23
Refinancing fee 4000
Total loan amount PV 57470.23
Term Nper 10
Rate Rate 4%
Annual payment PMT $7,085.56

Total payment under refinancing = 7085.56*10 = 70855.6

Payment without refinancing = 7264.89* 10 = 72648.9

Hence refinancing is a good option.

WORKINGS